# CogErgLLM: Exploring Large Language Model Systems Design Perspective Using Cognitive Ergonomics

### v1

@@ 1 Introduction
Though researchers are actively working in this area, current LLM system design often lacks comprehensive integration of cognitive ergonomics, resulting in systems that may not fully align with human cognitive capabilities and limitations. Secondly, there is insufficient focus on incorporating cognitive science methods to systematically identify and mitigate biases in LLM outputs. Additionally, existing LLM interfaces frequently fail to apply user-centered design principles consistently, leading to sub-optimal user experiences and cognitive overload. Lastly, LLMs often lack mechanisms to explain their decisions and outputs clearly, reducing user trust and transparency, while there’s also a noticeable gap in developing LLMs that can adapt to individual user preferences and learning styles over time, hindering their effectiveness and engagement (Eigner & Händler, 2024; Subramonyam et al., 2024; Liu et al., 2024; Le Guillou et al., 2022).
@@ 3.1 Methodology
The development methodology of the framework components, draws inspiration from theoretical foundations outlined in Sections 1 and A. These theoretical underpinnings of cognitive ergonomics serve as guiding principles for the design and definition of each framework element. We meticulously define and conceptualize the various components, informed by theories on human cognition and interaction with technology. Then, we transition to the implementation phase, where we translate the design aspects into tangible prototypes and undergo rigorous evaluation processes to assess their efficacy in enhancing user experience and system performance.
@@ 3.2.1 User-Centric Design
Personalization: Personalizing LLM interactions based on user profiles can significantly improve usability and satisfaction (Li et al., 2024). Cognitive ergonomics emphasizes the importance of designing systems that align with users’ mental models and preferences. By using data from user profiles, LLMs can offer tailored suggestions, responses, and content, making interactions more intuitive and reducing cognitive strain.
@@ 3.2.5 Trust and Transparency
Trust-building: Strategies for enhancing user trust in LLMs involve consistent, reliable performance, and the ability to demonstrate ethical and unbiased behavior. Incorporating user feedback and continuously improving the system based on that feedback also plays a crucial role in building and maintaining trust.
@@ 3.2.8 Personalization and Adaptation
Adaptive Learning: Integrating cognitive ergonomics principles into LLM design enables the development of systems that dynamically adapt to individual user preferences and learning styles. By leveraging cognitive science insights, such as theories of learning and memory, designers can create LLMs that personalize content delivery and interactions based on users’ behavior and performance, thereby enhancing learning outcomes and user satisfaction.
Personalized Recommendation: Cognitive ergonomics informs the implementation of personalized recommendation systems within LLMs, leveraging user data to offer tailored content and suggestions. By analyzing user interactions and preferences, designers can optimize content relevance and engagement, enhancing user experience and promoting continued usage of the LLMs (Lyu et al., 2024). Additionally, incorporating principles of cognitive load management ensures that recommendations are presented in a manner that optimizes and minimizes information overload.
@@ 5 Conclusion
Appendix A Background
A.1 Cognitive Ergonomics
A.2 Large Language Models and Cognition
Appendix B More Case Studies
Appendix C Discussion
Appendix D Challenges and Opportunities
D.1 Technical Challenges
D.2 Ethical Considerations
D.3 Future Opportunities

### v3

@@ 1 Introduction
Though researchers are actively working in this area, current LLM system design often lacks comprehensive integration of cognitive ergonomics, resulting in systems that may not fully align with human cognitive capabilities and limitations. Secondly, there is insufficient focus on incorporating cognitive science methods to systematically identify and mitigate biases in LLM outputs. Additionally, existing LLM interfaces frequently fail to apply user-centered design principles consistently, leading to sub-optimal user experiences and cognitive overload. Lastly, LLMs often lack mechanisms to explain their decisions and outputs clearly, reducing user trust and transparency, while there’s also a noticeable gap in developing LLMs that can adapt to individual user preferences and learning styles over time, hindering their effectiveness and engagement (Eigner & Händler, 2024; Subramonyam et al., 2024; Liu et al., 2024; Le Guillou et al., 2022; Wasi et al., 2024a).
@@
2 Background
2.1 Cognitive Ergonomics
2.2 Large Language Models and Cognition
@@ 4.1 Methodology
The development methodology of the framework components, draws inspiration from theoretical foundations outlined in Sections 1 and 2. These theoretical underpinnings of cognitive ergonomics serve as guiding principles for the design and definition of each framework element. We meticulously define and conceptualize the various components, informed by theories on human cognition and interaction with technology. Then, we transition to the implementation phase, where we translate the design aspects into tangible prototypes and undergo rigorous evaluation processes to assess their efficacy in enhancing user experience and system performance.
@@ 4.2.1 User-Centric Design
Personalization: Personalizing LLM interactions based on user profiles can significantly improve usability and satisfaction (Li et al., 2024; Wasi et al., 2024b). Cognitive ergonomics emphasizes the importance of designing systems that align with users’ mental models and preferences. By using data from user profiles, LLMs can offer tailored suggestions, responses, and content, making interactions more intuitive and reducing cognitive strain.
@@ 4.2.5 Trust and Transparency
Trust-building: Strategies for enhancing user trust in LLMs involve consistent, reliable performance, and the ability to demonstrate ethical and unbiased behavior (Wasi et al., 2024c). Incorporating user feedback and continuously improving the system based on that feedback also plays a crucial role in building and maintaining trust.
@@ 4.2.8 Personalization and Adaptation
Adaptive Learning: Integrating cognitive ergonomics principles into LLM design enables the development of systems that dynamically adapt to individual user preferences and learning styles. By leveraging cognitive science insights, such as theories of learning and memory, designers can create LLMs that personalize content delivery (Wasi et al., 2024b) and interactions based on users’ behavior and performance, thereby enhancing learning outcomes and user satisfaction.
Personalized Recommendation: Cognitive ergonomics informs the implementation of personalized recommendation systems within LLMs, leveraging user data to offer tailored content and suggestions. By analyzing user interactions and preferences, designers can optimize content relevance and engagement, enhancing user experience and promoting continued usage of the LLMs (Lyu et al., 2024; Wasi et al., 2024b). Additionally, incorporating principles of cognitive load management ensures that recommendations are presented in a manner that optimizes and minimizes information overload.
@@ 4.2.9 Continuous Evaluation and Improvement
4.3 Evaluation Metrics
Evaluation metrics provide a assessment of any implemented framework’s effectiveness and robustness. Usability metrics, cognitive load measurements, and user satisfaction ratings are used to evaluate the system’s ease of use, mental effort required, and overall user trust and acceptance.
1. Usability Metrics: Metrics such as System Usability Scale (SUS) scores (Brooke, 1995), task completion rates, and time-on-task measurements can assess the ease of use and efficiency of the LLM interfaces. High SUS scores and efficient task completion indicate a user-friendly design aligned with cognitive ergonomic principles.
2. Cognitive Load Metrics: Tools like the NASA Task Load Index (NASA-TLX) (nas, 1986) can measure the perceived cognitive workload of users interacting with the LLM. Lower cognitive load scores suggest that the information presentation and interaction design effectively reduce mental effort, aligning with the goals of cognitive ergonomics.
3. User Satisfaction and Trust: Surveys and feedback forms can gauge user satisfaction and trust in the system. Metrics such as Net Promoter Score (NPS) (Owen, 2018) and user satisfaction ratings provide insights into how well the system meets user expectations and builds trust through transparency and explainability.
@@ 5 Case Studies
6 Discussion
7 Challenges and Opportunities
7.1 Technical Challenges
7.2 Ethical Considerations
7.3 Future Opportunities